# Dynamic Approximate Maximum Matching in the Distributed Vertex Partition Model

### v2

@@ Abstract.
  • There is a randomized algorithm that succeeds with high probability in maintaining a $\frac{2}{3}$-approximate maximum matching and has a worst case update time of $\Omega\left(\frac{\ell\log n}{\sqrt{\beta k}}\right)$ rounds.
@@ 1. Introduction
Modern-day computing systems produce large amounts of data at an astonishing rate. These data sets often form complex relationships that are best represented as a graph consisting of vertices that are connected by edges. Performing efficient computations on such massive graphs is a fundamental challenge encountered in various applications, ranging from machine learning to protein interaction graphs and social network analysis. To cope with the scale of such massive graphs, which often exceeds the memory and bandwidth capabilities of a single machine, several iterative graph processing frameworks such as Google Pregel (Malewicz et al., 2010), and various open source projects such as, e.g., Apache Giraph (Ching et al., 2015), have been developed and are frequently used in industrial applications of graph analytics. In contrast to systems based on the MapReduce (Dean and Ghemawat, 2008) paradigm, these frameworks follow the vertex-centric programming model (Malewicz et al., 2010), where developers express their algorithms by “thinking like a vertex.” Intuitively speaking, this means that each vertex contains information about itself and its incident edges, and the computation is expressed at the level of a single vertex, which is very similar in spirit to traditional distributed graph algorithms (Peleg, 2000), and assumes that the input graph is vertex-partitioned among the physical machines forming the network. Moreover, some frameworks such as Giraph++ (Tian et al., 2013) and Blogel (Yan et al., 2014) expose a subgraph-centric programming interface in addition to the vertex-centric one, which provides greater flexibility by allowing algorithms that take into account the state of a subset of vertices as part of their state transition function.
1.1. Motivations and Open Questions
The general motivation of studying dynamic algorithms is to avoid computing anything from the scratch. Recall that computing a maximal matching in the $k\text{-}\mathsf{CLIQUE}_{\beta}$ model takes $\widetilde{O}(n/(k\beta))$ rounds (Li et al., 2024). Our goal is to maintain a maximal matching without 3-augmenting path using a small number of rounds for dynamic inputs.
In our paper, we do not design data structures to maintain the approximate maximum matching since it is not clear how to do that in distributed models, which leads to several open questions:
  • Can we save communication bits/rounds when maintaining memory? If so, how to design an efficient distributed data structure for dynamic problems?
  • What is the minimum size of memory required to maintain?
To answer these open questions in terms of lower bounds, one direction is to consider multi-party communication model with advice that is more challenging and exciting than two-party communication model with advice (see e.g., (Chattopadhyay et al., 2016)). For the upper bounds, it is really challenging to design an efficient distributed data structure for dynamic problems which is little explored in the literature.
@@ 1.2. The $k$-Clique Message Passing Model
When considering distributed graph algorithms in the big data setting, the size of the input graph typically dwarfs the number of available physical machines and the communication links between them. This is the underlying assumption of the $k$-clique message passing model (Li et al., 2024; Robinson, 2023) (denoted by $k\text{-}\mathsf{CLIQUE}_{\beta}$), where $k$ players $P_{1},\dots,P_{k}$ are interconnected using a clique topology of $\Theta(k^{2})$ point-to-point links with a per-link bandwidth of $\beta$ bits. In more detail, the $k\text{-}\mathsf{CLIQUE}_{\beta}$ model assumes that the input graph $G$ consists of $n\geqslant k$ vertices, $m$ edges, and $G$ is vertex-partitioned among the $k$ players in a roughly balanced way, such that each player obtains at most $O\left(\frac{n\log n}{k}\right)$ vertices. We assume a separation between the unit that stores (dynamic) input/output and the unit that performs computation which we call memory. All other data except input and output are stored in the memory.
@@
As mentioned, we focus on vertex-partitioning, which means that every vertex $u$ is assigned to some player $P$ together with its incident edges; consequently, every edge of the input graph is known to two players initially. Formally speaking, each vertex has a unique ID of $\Theta(\log n)$ bits, and “knowing” a vertex just means that this vertex’s ID is part of the local state of $P$. For each vertex $v$ assigned to $P$, we say that $P$ hosts vertex $v$, and we use $P_{u}$ to refer to the player hosting vertex $u$. The input of a player $P$ consists of the IDs of its hosted vertices and, for each of its vertices, $P$ knows the IDs of $v$’s neighbors as well as the player to which they were assigned to. In this work, we consider an adaptive adversary, who can choose any balanced vertex partitioning, which requires each player to hold at most $O\left(\frac{n}{k}\log n\right)$ vertices in total. [^1] Note that this is a main difference between the $k\text{-}\mathsf{CLIQUE}_{\beta}$ model and the $k$-machine model (Klauck et al., 2014), where the vertices are randomly partitioned.
We point out that, for $k=n$ players and $\beta=1$, our model is equivalent to the congested clique (Lotker et al., 2005) if each player obtains exactly one vertex. Analogously to the congested clique and the $\mathsf{CONGEST}$ model (Peleg, 2000) and in contrast to the popular Massively Parallel Computation (MPC) model (Karloff et al., 2010), we do not impose a restriction on the local memory of the players in the $k\text{-}\mathsf{CLIQUE}_{\beta}$ model. That said, our results for single updates hold even for memoryless dynamic algorithms that do not use any additional memory apart from storing the input and output.
@@ 1.3. Dynamic Updates
To address the reality of evolving data sets in the real-world, we assume that the initial graph consists of $n$ vertices without any edges, and model dynamical changes by considering a sequence of updates to the edges of the graph. There are two important models in this setting: single-updates where each update is just the addition or the deletions of a single edge, and batch-dynamic updates, where each update consists of a batch of at most $\ell$ edge deletions/insertions. [^2] Batch-dynamic algorithms have attracted significant attention recently, see (Acar et al., 2019, 2011; Anderson, 2023; Dhulipala et al., 2020, 2021; Liu et al., 2022; Nowicki and Onak, 2021; Gilbert and Li Er Lu, 2020; Li et al., 2024). In either case, the adaptive adversary may observe the local memory of the players, which includes the currently computed output, before choosing the next update. Since we are assuming a distributed algorithm, each player only needs to output the part of the solution relevant to its hosted vertices. That is, for computing a matching $M$, each player $P$ needs to output every edge $\{u,v\}\in M$, for which it hosts an endpoint.
@@
Approximate Maximum Matching. In this work, we focus on computing a matching, which is one of the most fundamental problems in computer science. A matching in a graph $G$ is a set of edges without common endpoints, and we say that a matching $M$ is a maximum (cardinality) matching if no matching in $G$ is larger than $M$. A matching $M$ is an $\alpha$-approximate matching if $|M|\geqslant\alpha|\textsc{mcm}(G)|$ where $\textsc{mcm}(G)$ is any matching in $G$. It is well-known that if there is no augmenting paths of length 2k12k-1 or less, the obtained matching is at least a $\frac{k}{k+1}$-approximation, see (Hopcroft and Karp, 1973), which we leverage in our algorithms.
1.4. Our Contributions
@@
Next, we turn our attention to the batch-dynamic model, where the adversary can perform up to $\ell$ updates instantaneously. It is not clear how to handle just one single edge deletion efficiently. Therefore, we focus on the incremental setting, where the adversary can only insert batches of edges. Our main technical contribution is an algorithm that handles batches of edge insertions efficiently:
@@
Our algorithm for Result 2 makes use of certain structural properties of the graph for resolving 3-augmenting paths, in addition to employing the maximal matching algorithm of (Li et al., 2024) as a subroutine. In more detail, we first identify all 3-augmenting paths that traverse two newly inserted edges, as well as “triangles” that lead to 3-augmenting paths, and resolve all of these locally by sending the induced subgraph to a player. Then, we can focus on 3-augmenting paths that traverse only a single inserted edge, and we point out that there can be a large number of them, which prohibits local processing. Instead, we consider a much smaller subgraph, called $G_{*}^{\prime}$, which includes the edges of a certain maximal matching on a bipartite graph. We show that it is sufficient to break up 3-augmenting paths in $G_{*}^{\prime}$ to avoid any 3-augmenting paths in the entire graph.
To visualize Result 2 for large-scale data, we can set the bandwidth of each machine [^3] Note that in another large-scale distributed model, i.e., the MPC model, the memory of each machine can be as large as $O(n)$. to be $O(n)$ by adjusting $\beta$. And then we can handle $O(\sqrt{n})$ incremental updates using $O(\log n)$ rounds, which is more interesting and challenging than handling one update for dynamic $2/3$-approximate maximum matching in the MPC model  (Italiano et al., 2019).
@@ 1.5. Additional Related Work
Most closely related to our work is  (Li et al., 2024), who study dynamic maximal matching in the $k\text{-}\mathsf{CLIQUE}_{\beta}$ model. They give lower bounds and corresponding algorithms for both oblivious as well as adaptive adversaries. Notice that a maximal matching is a local problem, in contrast to the more challenging “global” problem of computing an (approximate) maximum matching. Dynamic $2/3$-approximate maximum matching was also stuided in the MPC model. The paper (Italiano et al., 2019) gave a constant-round dynamic algorithm for $2/3$-approximate maximum matching for single updates, instead of batches of updates. Their idea is to maintain a maximal matching without 3-augmenting path, which is similar to the idea in this paper. We should notice that in this paper, we handle batches of updates, more challenging and complicated. For approximate maximum matching in the MPC model, (Ghaffari et al., 2018) proposes a $(1+\epsilon)$-approximate maximum matching in $O(\log\log n)$ rounds with $\widetilde{O}(n)$ memory per machine. By simulating their algorithm in the $k\text{-}\mathsf{CLIQUE}_{\beta}$ model, we can obtain a $(1+\epsilon)$-approximate maximum matching in roughly $O(\lceil\frac{m}{k^{2}\beta}\rceil+\frac{n}{k\beta})$ rounds if each vertex of the input graph and each of its incident edges assigned to machines are chosen independently and uniformly at random.
@@
We next discuss related work in the centralized settings. There is a survey on recent advances in fully dynamic algorithms (Hanauer et al., 2022). Edge degree constrained subgraphs (EDCS) (Bernstein and Stein, 2015) is an elegant and popular data structure to maintain matching for dynamic changes in centralized settings. To the best of our knowledge, the first incremental maximum matching is given by Gupta (Gupta, 2014). Subsequently, (Grandoni et al., 2019) shows a deterministic algorithm that maintains a $(1+\varepsilon)$-approximate matching with constant amortized update time per insertion. The techniques of (Grandoni et al., 2019) are based on augmenting paths elimination. Recently, (Blikstad and Kiss, 2023) further improved the update time complexity for maintaining an incremental $(1-\varepsilon)$-approximate matching to be $O(\operatorname{\mathrm{poly}}(1/\varepsilon))$ by weighted EDCS. On the lower bound side, it is known that maintaining an exact matching in the incremental/decremental settings require $\Omega(n^{0.333-o(1)})$ update time assuming the 3-SUM conjecture (Kopelowitz et al., 2016). Even for bipartite graphs, the lower bound for maintaining a maximum matching in the incremental graph needs $\Omega(n^{1-\varepsilon})$ (Dahlgaard, 2016) under the OMv conjecture.
@@
From the above description of the progress of dynamic (approximate) maximum matching in centralized settings, we can see that EDCS is a crucial tool for dynamic matching. Unfortunately, in most distributed settings, especially for models with small bandwidth, it is hard to maintain an EDCS because of the restricted communication bandwidth. For maintaining a $\frac{2}{3}$-approximate maximum matching, a natural idea is to keep that the graph has no augmenting paths of length three. In the MPC model (where the total bandwidth can be $\widetilde{O}(m)$ where $m$ is the number of edges of the input graph), the algorithm of (Italiano et al., 2019) can update a $\frac{2}{3}$-approximate cardinality maximum matching in constant rounds when considering a single update. Notice that in clique networks with small bandwidth, it would be expensive to achieve this goal if we find such augmenting paths by searching neighborhoods. While an alternative way is to maintain a 2-hop neighborhood, doing so is prohibitively expensive in the distributed setting.
@@
Approximate maximum matching has been studied extensively the streaming setting. It is straightforward to transform single-pass semi-streaming algorithms to incremental algorithms in the $k\text{-}\mathsf{CLIQUE}_{\beta}$ model. Unfortunately, the approximation ratio obtained from the transformation cannot be as good as $2/3$, as (Kapralov, 2013) shows that there is no single-pass semi-streaming algorithm for achieving $1-1/e\approx 0.63<2/3$ approximation to maximum matching. On the other hand, multi-pass streaming algorithms can obtain $(1-\epsilon)$-approximate maximum matching (see the survey (Assadi, 2023)). An interesting direction for future exploration is to transform multiple-pass streaming algorithms to incremental algorithms in the distributed settings.
@@
When we consider edge insertions/deletions in streams, it is corresponding to dynamic streams (or turnstile streams [^4] They are nearly the same thing in the literature.). Dynamic streaming algorithms can be considered as dynamic algorithms with limited space. Due to this property, there are strong relationships between dynamic streaming algorithms and dynamic algorithms in the distributed models. The paper (Ahn et al., 2012) initiates the study of dynamic graph streaming algorithms. After that, (Assadi et al., 2016) shows that the space complexity of single-pass turnstile streaming algorithms for $n^{\epsilon}$-approximate maximum matching is $\Omega(n^{2-3\epsilon})$, and a similar lower bound for bipartite matching with expected approximation $n^{\epsilon}$ is also given in (Konrad, 2015). For random streams, i.e., edges arrive in a random order,  (Farhadi et al., 2020) proposes a single-pass semi-streaming algorithm that finds approximately 0.545-approximation of the maximum matching in a general graph.

### v3

# Approximate Maximum Matching in the Distributed Vertex Partition Model under Dynamic Graph Updates

@@ Abstract.
  • There is a randomized algorithm that succeeds with high probability in maintaining a $\frac{2}{3}$-approximate maximum matching and has a worst case update time of $O\left(\left\lceil\frac{\ell\log n}{\sqrt{\beta k}}\right\rceil\right)$ rounds.
@@ 1. Introduction
Modern-day computing systems produce large amounts of data at an astonishing rate. These data sets often form complex relationships that are best represented as a graph consisting of vertices that are connected by edges. Performing efficient computations on such massive graphs is a fundamental challenge encountered in various applications, ranging from machine learning to protein interaction graphs and social network analysis. To cope with the scale of such massive graphs, which often exceeds the memory and bandwidth capabilities of a single machine, several iterative graph processing frameworks such as Google Pregel (pregel), and various open source projects such as, e.g., Apache Giraph (facebookgiraph), have been developed and are frequently used in industrial applications of graph analytics. In contrast to systems based on the MapReduce (dean2008mapreduce) paradigm, these frameworks follow the vertex-centric programming model (pregel), where developers express their algorithms by “thinking like a vertex.” Intuitively speaking, this means that each vertex contains information about itself and its incident edges, and the computation is expressed at the level of a single vertex, which is very similar in spirit to traditional distributed graph algorithms (peleg_book), and assumes that the input graph is vertex-partitioned among the physical machines forming the network. Moreover, some frameworks such as Giraph++ (tian2013think) and Blogel (yan2014blogel) expose a subgraph-centric programming interface in addition to the vertex-centric one, which provides greater flexibility by allowing algorithms that take into account the state of a subset of vertices as part of their state transition function.
@@ 1.1. The $k$-Clique Message Passing Model
When considering distributed graph algorithms in the big data setting, the size of the input graph typically dwarfs the number of available physical machines and the communication links between them. This is the underlying assumption of the $k$-clique message passing model (li2024dynamic; amg) (denoted by $k\text{-}\mathsf{CLIQUE}_{\beta}$), where $k$ players $P_{1},\dots,P_{k}$ are interconnected using a clique topology of $\Theta(k^{2})$ point-to-point links with a per-link bandwidth of $\beta$ bits. In more detail, the $k\text{-}\mathsf{CLIQUE}_{\beta}$ model assumes that the input graph $G$ consists of $n\geqslant k$ vertices, $m$ edges, and $G$ is vertex-partitioned among the $k$ players in a roughly balanced way, such that each player obtains at most $O\left(\frac{n\log n}{k}\right)$ vertices. We assume a separation between the unit that stores (dynamic) input/output and the unit that performs computation which we call memory. All other data except input and output are stored in the memory.
@@
As mentioned, we focus on vertex-partitioning, which means that every vertex $u$ is assigned to some player $P$ together with its incident edges; consequently, every edge of the input graph is known to two players initially. Formally speaking, each vertex has a unique ID of $\Theta(\log n)$ bits, and “knowing” a vertex just means that this vertex’s ID is part of the local state of $P$. For each vertex $v$ assigned to $P$, we say that $P$ hosts vertex $v$, and we use $P_{u}$ to refer to the player hosting vertex $u$. The input of a player $P$ consists of the IDs of its hosted vertices and, for each of its vertices, $P$ knows the IDs of $v$’s neighbors as well as the player to which they were assigned to. In this work, we consider an adaptive adversary, who can choose any balanced vertex partitioning, which requires each player to hold at most $O\left(\frac{n}{k}\log n\right)$ vertices in total. [^1] Note that this is a main difference between the $k\text{-}\mathsf{CLIQUE}_{\beta}$ model and the $k$-machine model (klauck2014distributed), where the vertices are randomly partitioned.
We point out that, for $k=n$ players and $\beta=1$, our model is equivalent to the congested clique (lotker) if each player obtains exactly one vertex. Analogously to the congested clique and the $\mathsf{CONGEST}$ model (peleg_book) and in contrast to the popular Massively Parallel Computation (MPC) model (mpc), we do not impose a restriction on the local memory of the players in the $k\text{-}\mathsf{CLIQUE}_{\beta}$ model. That said, our results for single updates hold even for memoryless dynamic algorithms that do not use any additional memory apart from storing the input and output.
@@ 1.2. Dynamic Updates
To address the reality of evolving data sets in the real-world, we assume that the initial graph consists of $n$ vertices without any edges, and model dynamical changes by considering a sequence of updates to the edges of the graph. There are two important models in this setting: single-updates where each update is just the addition or the deletions of a single edge, and batch-dynamic updates, where each update consists of a batch of at most $\ell$ edge deletions/insertions. [^2] Batch-dynamic algorithms have attracted significant attention recently, see (acar2019parallel; acar2011parallelism; anderson2023parallel; dhulipala2020parallel; dhulipala2021parallel; liu2022parallel; nowicki2021dynamic; gilbert2020fast; li2024dynamic). In either case, the adaptive adversary may observe the local memory of the players, which includes the currently computed output, before choosing the next update. Since we are assuming a distributed algorithm, each player only needs to output the part of the solution relevant to its hosted vertices. That is, for computing a matching $M$, each player $P$ needs to output every edge $\{u,v\}\in M$, for which it hosts an endpoint.
@@
Approximate Maximum Matching. In this work, we focus on computing a matching, which is one of the most fundamental problems in computer science. A matching in a graph $G$ is a set of edges without common endpoints, and we say that a matching $M$ is a maximum (cardinality) matching if no matching in $G$ is larger than $M$. A matching $M$ is an $\alpha$-approximate matching if $|M|\geqslant\alpha|\textsc{mcm}(G)|$ where $|\textsc{mcm}(G)|$ is the cardinality of the maximum matching. It is well-known that the obtained matching is at least a $\frac{k}{k+1}$-approximation, if there is no augmenting paths of length 2k12k-1 or less, see (hopcroft1973n), which we leverage in our algorithms:
Lemma 1 ((hopcroft1973n)).
1.3. Contributions and Technical Challenges
@@
For obtaining Result 1, we show an edge-insertion can be dealt with by simply coordinating between the players that hold a vertex adjacent to the new edge and checking whether adding this edge to the matching induces a 3-augmenting path in the graph.
Handling an edge-deletion is more challenging, however, since a deleted matched edge may cause many nodes to become “free” (i.e., unmatched), which in turn may result in 3-augmenting paths that are difficult to find in sufficiently dense graphs. More concretely, suppose that the matched edge $\{u,v\}$ was deleted and $u$ (1) has high node degree and (2) no free neighbor. We show that we can sample some neighbor $w$ of $u$ that itself has low degree, remove $w$’s edge from the matching and instead add the edge $\{u,w\}$. As this operation may result in new 3-augmenting paths, we leverage the fact that $w$ has low degree, which allows us use an information dissemination routine of (li2024dynamic) to find and resolve such a 3-augmenting path if it indeed exists.
Next, we turn our attention to the batch-dynamic model, where the adversary can perform up to $\ell$ updates instantaneously. As described above, handling even just a single edge-deletion requires a non-negligible amount of communication, and thus we exclusively focus on the batch-incremental setting, where the adversary can insert batches of edges. Our main technical contribution is an algorithm that handles batches of edge-insertions efficiently:
@@
Our algorithm for Result 2 makes use of certain structural properties of the graph for resolving 3-augmenting paths, in addition to employing the maximal matching algorithm of (li2024dynamic) as a subroutine. In more detail, we first identify all 3-augmenting paths that traverse two newly inserted edges, as well as “triangles” that lead to 3-augmenting paths, and resolve all of these locally by sending the induced subgraph to a player. Then, we can focus on 3-augmenting paths that traverse only a single inserted edge, and we point out that there can be a large number of them, which prohibits local processing. Instead, we consider a much smaller subgraph, called $G_{*}^{\prime}$, which includes the edges of a certain maximal matching on a bipartite graph. We show that it is sufficient to break up 3-augmenting paths in $G_{*}^{\prime}$ to avoid any 3-augmenting paths in the entire graph.
To give a concrete instance of the upper bound stated in Result 2 for the large-scale data setting, we can set the link bandwidth $\beta$ to be $\Theta(n)$. Under this assumption, we can handle batches of $O(\sqrt{n})$ incremental updates in just $O(\log n)$ rounds.
@@ 1.4. Additional Related Work
Most closely related to our work is  (li2024dynamic), who study dynamic maximal matching in the $k\text{-}\mathsf{CLIQUE}_{\beta}$ model. They give lower bounds and corresponding algorithms for both oblivious as well as adaptive adversaries. Notice that a maximal matching is a local problem, in contrast to the more challenging “global” problem of computing an (approximate) maximum matching. Dynamic $2/3$-approximate maximum matching was also stuided in the MPC model. The paper (italiano2019dynamic) gave a constant-round dynamic algorithm for $2/3$-approximate maximum matching for single updates, instead of batches of updates. Their idea is to maintain a maximal matching without 3-augmenting path, which is similar to the idea in this paper. We should notice that in this paper, we handle batches of updates, more challenging and complicated. For approximate maximum matching in the MPC model, (ghaffari2018improved) proposes a $(1+\epsilon)$-approximate maximum matching in $O(\log\log n)$ rounds with $\widetilde{O}(n)$ memory per machine. By simulating their algorithm in the $k\text{-}\mathsf{CLIQUE}_{\beta}$ model, we can obtain a $(1+\epsilon)$-approximate maximum matching in roughly $O(\lceil\frac{m}{k^{2}\beta}\rceil+\frac{n}{k\beta})$ rounds if each vertex of the input graph and each of its incident edges assigned to machines are chosen independently and uniformly at random.
@@
We next discuss related work in the centralized settings. There is a survey on recent advances in fully dynamic algorithms (hanauer2022recent). Edge degree constrained subgraphs (EDCS) (bernstein2015fully) is an elegant and popular data structure to maintain matching for dynamic changes in centralized settings. To the best of our knowledge, the first incremental maximum matching is given by Gupta (gupta2014maintaining). Subsequently, (grandoni2019) shows a deterministic algorithm that maintains a $(1+\varepsilon)$-approximate matching with constant amortized update time per insertion. The techniques of (grandoni2019) are based on augmenting paths elimination. Recently, (blikstad2023incremental) further improved the update time complexity for maintaining an incremental $(1-\varepsilon)$-approximate matching to be $O(\operatorname{\mathrm{poly}}(1/\varepsilon))$ by weighted EDCS. On the lower bound side, it is known that maintaining an exact matching in the incremental/decremental settings require $\Omega(n^{0.333-o(1)})$ update time assuming the 3-SUM conjecture (kopelowitz2016higher). Even for bipartite graphs, the lower bound for maintaining a maximum matching in the incremental graph needs $\Omega(n^{1-\varepsilon})$ (dahlgaard2016hardness) under the OMv conjecture.
@@
From the above description of the progress of dynamic (approximate) maximum matching in centralized settings, we can see that EDCS is a crucial tool for dynamic matching. Unfortunately, in most distributed settings, especially for models with small bandwidth, it is hard to maintain an EDCS because of the restricted communication bandwidth. For maintaining a $\frac{2}{3}$-approximate maximum matching, a natural idea is to keep that the graph has no augmenting paths of length three. In the MPC model (where the total bandwidth can be $\widetilde{O}(m)$ where $m$ is the number of edges of the input graph), the algorithm of (italiano2019dynamic) can update a $\frac{2}{3}$-approximate cardinality maximum matching in constant rounds when considering a single update. Notice that in clique networks with small bandwidth, it would be expensive to achieve this goal if we find such augmenting paths by searching neighborhoods. While an alternative way is to maintain a 2-hop neighborhood, doing so is prohibitively expensive in the distributed setting.
@@
Approximate maximum matching has been studied extensively the streaming setting. It is straightforward to transform single-pass semi-streaming algorithms to incremental algorithms in the $k\text{-}\mathsf{CLIQUE}_{\beta}$ model. Unfortunately, the approximation ratio obtained from the transformation cannot be as good as $2/3$, as (kapralov2013better) shows that there is no single-pass semi-streaming algorithm for achieving $1-1/e\approx 0.63<2/3$ approximation to maximum matching. On the other hand, multi-pass streaming algorithms can obtain $(1-\epsilon)$-approximate maximum matching (see the survey (assadi2023recent)). An interesting direction for future exploration is to transform multiple-pass streaming algorithms to incremental algorithms in the distributed settings.
@@
When we consider edge insertions/deletions in streams, it is corresponding to dynamic streams (or turnstile streams [^3] They are nearly the same thing in the literature.). Dynamic streaming algorithms can be considered as dynamic algorithms with limited space. Due to this property, there are strong relationships between dynamic streaming algorithms and dynamic algorithms in the distributed models. The paper (ahn2012analyzing) initiates the study of dynamic graph streaming algorithms. After that, (assadi2016maximum) shows that the space complexity of single-pass turnstile streaming algorithms for $n^{\epsilon}$-approximate maximum matching is $\Omega(n^{2-3\epsilon})$, and a similar lower bound for bipartite matching with expected approximation $n^{\epsilon}$ is also given in (konrad2015maximum). For random streams, i.e., edges arrive in a random order,  (farhadi2020approximate) proposes a single-pass semi-streaming algorithm that finds approximately 0.545-approximation of the maximum matching in a general graph.